Analytic Number Theory and Surrounding Areas

Date : 18-22 October, 2004 Place : Research Institute of Mathematical Science (RIMS), Kyoto, JAPAN Organizer : Masaaki Amou (Gunma Univ.)

Program

Monday 18 October

13:25〜13:30 Opening
13:30〜14:20 Shigeki Akiyama (Niigata Univ.)

3/2 number system and Mahler’s problem

14:30〜15:00 Taka-aki Tanaka (Keio Univ.)

Algebraic independence of the values of a certain function at distinct algebraic points

15:10〜15:40 Yohei Tachiya (Keio Univ.)

Transcendence of certain infinite products

16:00〜16:30 Masahiro Asano (Nagoya Univ.)

A generalization of the reciprocity law of multiple Dedekind sums

16:40〜17:20 Masatoshi Suzuki (Nagoya Univ.)

The Riemann Hypothesis for certain integrals of Eisenstein sereis Tuesday 19 October

9:00〜9:50 Jianya Liu (Shandong Univ.)
     Bounds for automorphic L-functions
10:00〜10:50 Kohji Matsumoto (Nagoya Univ.)

The discrete universality of automorphic L-functions

11:00〜11:50 Winfried Kohnen (Univ. Heidelberg)

On q-product expansions of certain modular functions

13:30〜14:00 Yuichi Kamiya

The Beurling spectrum of certain zeta-functions

14:10〜14:40 Yumiko Ichihara (Hiroshima Univ.)

On the Siegel-Tatuzawa theorem for a class of L-functions

15:00〜15:50 Clemens Heuberger (Technische Univ. Graz)
      Parametrized Families of Relative Thue Equations
16:00〜16:50 Ryotaro Okazaki (Doshisya Univ.)

Simultaneous Pell Equation of Indefinite Signature

17:00〜17:50 Noriko Hirata-Kohno (Nihon Univ.)

S-unit Equations and Integer Solutions to

Exponential Diophantine Equations

Wednesday 20 October

9:30〜10:30 Alf van der Poorten (Macquarie Univ.)
      Elliptic Sequences and Continued Fractions
10:45〜11:45 Tanguy Rivoal (Univ. Caen)
      Hypergeometry and the Riemann zeta function

Afternoon  Free Time

Thursday 21 October

9:00〜10:00 Imre Kátai (Eotvos Lorand Univ.)

Distribution of arithmetical functions on some subsets of integers

10:10〜11:00 Francesco Pappalardi (Univ. degli Studi Roma Tre)
      Values of the Carmichael function versus values of the

Euler function

11:10〜11:40 Temenoujka Peneva (Plovdiv Univ., Tsukuba Univ.)

Hiroshi Mikawa (Tsukuba Univ.) A variant of Bombieri-Vinogradov theorem

13:30〜14:00 Koichi Kawada (Iwate Univ.)

On treatment of minor arc integrals in applications of the circle method

14:10〜15:00 Alain Lasjaunias (Univ. Bordeaux I)

Diophantine Approximation in Positive Characteristic

15:10〜15:40 Takao Komatsu (Hirosaki Univ.)

Rational approximations to Tasoev type continued fractions

15:50〜16:20 Rie Natsui (Keio Univ.)

On the arithmetic distributions of simultaneous approximation convergents arising from Jacobi-Perron algorithm

16:30〜17:20 Jun-ichi Tamura

Hankel determinants and substitutions

18:00〜20:00 Reception at Kyodai Kaikan

Friday 22 October

9:30〜10:20 Kalyan Chakraborty (Harish-Chandra Research Inst.)
     On the divisibility of class numbers of real quadratic fields
10:30〜11:30 Junjiro Noguchi (Tokyo Univ.)

Recent results on Analogue in Nevanlinna Theory and Diophantine Approximation

13:30〜14:00 Masami Fujimori (Kanagawa Inst. Tech.)

On the defining group of a category of systems of linear inequalities

14:10〜14:40 Masami Yoshimoto (Nagoya Univ.)

On the functional equation of the Barnes multiple zeta-function

14:50〜15:20 Masanori Katsurada (Keio Univ.)

Complete asymptotic expansions associated with Epstein zeta-functions